Use the given information to find the Upper P​-value. The test statistic in a​ two-tailed test is z= 1.67. Use the Cumulative Probabilities Table
Answer #1

Solution :

Given that,

This is the two tailed test ,

z = 1.67

Using standard normal table,

P(z > 1.67) = 1 - P(z < 1.67) = 1 - 0.9525 = 0.0475

P-value = 2 * P(z > 1.67) = 2 * 0.0475 = 0.095

P-value = 0.095